Income Elasticity of Demand

Income elasticity of demand formula: EY= Q2 - Q1/(Q2 + Q1)/2 divided by Y2 - Y1/(Y2 + Y1)/2 (remember that income elasticity measures how a change in income effects the demand for a good)

Solve the following problem: Bob makes $30,000 per year and eats at local eateries two nights a week. When he received a raise, and now makes $40,000 per year, he increased his dining out to five nights per week. Provide your answer, and indicate if your number showed elasticity, unitary elasticity, or was inelastic.
Answer #1

Here, Y1 = 30000, Y2 = 40000, Q1 = 2, Q2 = 5

EY = (Q2 - Q1) / (Y2 - Y1) * (Y1 + Y2) / (Q1 + Q2)

= (5 - 2) / (40000 - 30000) * (30000 + 40000) / (2 + 5)

= (3 / 10000) * (70000 / 7)

= 3

Since EY is greater than 1, it is elastic.